OpenStudy (anonymous):

find the center (h,k) and the radius, r, of the circle. 4x^2-4x+4y^2-9y-3=0

OpenStudy (anonymous):

4x^2-4x +4y^2-9y=3 Complete the square

OpenStudy (anonymous):

divide through by 4

OpenStudy (anonymous):

x^2 -x +y^2 -(9/4)y = 3/4 (x-(1/2) )^2 + (y - (9/2) )^2 = (3/4) + (1/4) + (9/4)

OpenStudy (anonymous):

(x-(1/2) ) ^2 + (y- ( 9/2) ) ^2 = 4 centre ( (1/2) , (9/2) ) radius 2